# More is Different: Multi-Axion Dynamics Changes Topological Defect Evolution

### v1

@@ I Introduction
The results reveal that the structure within the core of axion strings at low energies consists of an intricate assembly of strings and DWs, termed a “string bundle” Higaki:2016jjh . This string bundle contains $N_{\rm axion}$ types of strings, with an exponential hierarchy required among their numbers. Consequently, forming such a string bundle from individual strings that evolve following the scaling solution would take an exponentially long time. As a result, in realistic cosmic evolution, it becomes impossible to form a string bundle, and the UV physics that should be confined to the axion string core spreads throughout the universe. Specifically, DWs with large tension are considered to extend across the cosmic scales. As stressed in Ref. Higaki:2016jjh , the decay of these high-tension DWs can generate extremely strong gravitational waves, which can be probed by observations such as pulsar timing arrays. Interestingly, such axion DWs in multi-axion systems potentially explain the recent hint for stochastic gravitational waves found by the pulsar timing arrays such as NANOGrav NANOGrav:2023gor ; Antoniadis:2023ott ; Reardon:2023gzh ; Xu:2023wog . See Refs. Kitajima:2023cek ; Ferreira:2024eru ; Dankovsky:2024zvs for the domain-wall explanation of the NANOGrav signal.
@@
Finally, we comment on the relevant literature on multiple axions, and mixed pre- and post-inflationary initial conditions. The topological defects in multiple axion models were studied in Refs. Higaki:2016jjh ; Long:2018nsl ; Hiramatsu:2019tua ; Lee:2024xjb . In our recent work Lee:2024xjb , we studied a two-axion system with mixed post- and pre-inflationary initial conditions. We showed that the string-wall network associated with one axion with the post-inflationary initial condition can induce a domain-wall-like structure for the other axion with the pre-inflationary initial condition. As we will show later, these “induced domain walls” play a crucial role in multi-axion systems, particularly in understanding the stability of the string-wall network. In this paper, when we consider pre-inflationary conditions, we focus on scenarios where the corresponding axion strings have been sufficiently diluted by inflation to be effectively absent from the observable universe. It is worth noting, however, that there are scenarios considering a very sparse distribution of strings. Such scenarios can arise if SSB occurs during or near the end of inflation Lazarides:1984pq ; Shafi:1984tt ; Vishniac:1986sk ; Yokoyama:1988zza , or if late-time inflation, such as thermal inflation Yamamoto:1985rd ; Lazarides:1985ja ; Lyth:1995hj ; Lyth:1995ka , is considered. Recently, Ref. Bao:2024bws examined a system of two mixing axions, applying post-inflationary conditions to one and this highly diluted string distribution to the other, focusing on the generation of gravitational waves.
@@ III.1.1 Post-post scenario
Figure 2: The snapshots of the energy density at the final time $t=300/\sqrt{\lambda}\eta$ for $(n_{1},n_{2})=(1,2)$ (upper panel) and $(3,2)$ (lower panel) are shown. The energy density is normalized by $\lambda\eta^{4}$. In the upper panel, several string bundles, each composed of three strings, are formed. In contrast, a string-wall network is formed in the lower panel. In both panels, one side of the box is about the size of two Hubble horizons.
@@ III.2.2 Post-post scenario
On the other hand, if $D=0$, the $U(1)_{L}$ symmetry remains unbroken, and there is no DW attached to the string bundle. Thus, the string bundles without DWs simply follow the scaling solution. Let us mention here a potential difference from the simple single-axion case. Since the string bundles have an internal structure, it is nontrivial whether they evolve exactly in the same way as ordinary axion strings. For instance, if a string bundle forms a twisted loop, its contraction will stop at a certain radius due to the existence of multiple strings inside the bundle. Then, such twisted string loops are stable and may contribute to dark matter. A detailed investigation of the formation of such loops and other exotic configurations of topological defects requires dedicated numerical simulations, which warrant further investigation.
@@ III.4 Summary of classification
$\geq 1$ $1$ 00 $\theta_{L}$ string bundles $\theta_{L}$ string bundles Fig. 2 -
@@ IV Cosmological implications
Alternatively, the string-wall network can collapse due to a transient bias that arises before the axions settle to the global minimum of the potential. This transient bias may either directly collapse the network or indirectly destabilize the network by introducing a population bias. As a result, the timescale of the string-wall network collapse can vary over a wide range.
@@
Thus, the key features of the QCD axion model with multiple axions are the general existence of high-tension DW networks and the generation of gravitational waves accompanying their collapse. This is precisely what was pointed out in the context of the clockwork QCD axion in Ref. Higaki:2014pja .555 Ref. Higaki:2014pja considered post-inflationary conditions and stated that if $N_{\rm DW}=1$, DWs would generally collapse if there were $N_{\rm axion}$ potential terms for $N_{\rm axion}$ axions. However, we have pointed out that this part is not correct, as the existence of induced walls was overlooked. Here, we have seen that this holds true even in the case of multiple axions, particularly for two axions.
@@ V Summary
We introduced two hierarchical axion potential terms and examined the evolution of topological defects based on their DW numbers. In the post-post scenario, configurations of string bundles with heavy axions confined within them exist but form effectively only under specific initial conditions and DW numbers. More generally, we observe stable networks with high-tension DWs caused by heavy axions. This challenges the common assumption that heavy axions can be integrated out to simplify the system to a single-axion model, especially in the presence of cosmic strings. In the pre-post scenario, the formation of such string bundles is impossible due to one set of strings being diluted away by inflation, inevitably leading to DW networks. The ubiquity of these DW networks in multi-PQ scalar scenarios is a key argument of this paper, generalizing the observation made by Higaki:2014pja in the context of clockwork axions.

### v2

@@ I Introduction
The results reveal that the structure within the core of axion strings at low energies consists of an intricate assembly of strings and DWs, termed a “string bundle” Higaki:2016jjh . This string bundle contains $N_{\rm axion}$ types of strings, with an exponential hierarchy required among their numbers. Consequently, forming such a string bundle from individual strings that evolve following the scaling solution would take an exponentially long time. As a result, in realistic cosmic evolution, it becomes impossible to form a string bundle, and the UV physics that should be confined to the axion string core spreads throughout the universe. Specifically, DWs with large tension are considered to extend across the cosmic scales. As stressed in Ref. Higaki:2016jjh , the decay of these high-tension DWs can generate extremely strong gravitational waves, which can be probed by observations such as pulsar timing arrays. Interestingly, such axion DWs in multi-axion systems potentially explain the recent hint for stochastic gravitational waves found by the pulsar timing arrays such as NANOGrav NANOGrav:2023gor ; Antoniadis:2023ott ; Reardon:2023gzh ; Xu:2023wog . See Refs. Kitajima:2023cek ; Ferreira:2024eru ; Dankovsky:2024zvs ; Blasi:2023sej for the domain-wall explanation of the NANOGrav signal.
@@
Finally, we comment on the relevant literature on multiple axions, and mixed pre- and post-inflationary initial conditions. The topological defects in multiple axion models were studied in Refs. Higaki:2016jjh ; Long:2018nsl ; Hiramatsu:2019tua ; Eto:2023aqr ; Lee:2024xjb . In our recent work Lee:2024xjb , we studied a two-axion system with mixed post- and pre-inflationary initial conditions. We showed that the string-wall network associated with one axion with the post-inflationary initial condition can induce a domain-wall-like structure for the other axion with the pre-inflationary initial condition. As we will show later, these “induced domain walls” play a crucial role in multi-axion systems, particularly in understanding the stability of the string-wall network. In this paper, when we consider pre-inflationary conditions, we focus on scenarios where the corresponding axion strings have been sufficiently diluted by inflation to be effectively absent from the observable universe. It is worth noting, however, that there are scenarios considering a very sparse distribution of strings. Such scenarios can arise if SSB occurs during or near the end of inflation Lazarides:1984pq ; Shafi:1984tt ; Vishniac:1986sk ; Yokoyama:1988zza , or if late-time inflation, such as thermal inflation Yamamoto:1985rd ; Lazarides:1985ja ; Lyth:1995hj ; Lyth:1995ka , is considered. Recently, Ref. Bao:2024bws examined a system of two mixing axions, applying post-inflationary conditions to one and this highly diluted string distribution to the other, focusing on the generation of gravitational waves. In the context of the string axiverse, Ref. Benabou:2023npn studied axion strings and DWs in a two-axion system, similar to the pre-post scenario we will discuss later. Their work, however, did not address the induced DWs. We will show that the induced DWs play a crucial role in determining the network’s stability.
@@ III.1.1 Post-post scenario
We have considered a two-axion system without a hierarchy in the decay constants to discuss the formation of string bundles. Note that if a hierarchy between the decay constants exists, the DWs take longer to shrink, as we show here. For the case where $|n_{1}|=1$ and $|n_{2}|=\mathcal{O}(1)$ but not equal to 1 (or vice versa), the DWs bounded by strings contract soon after the DW tension force becomes comparable to the string tension force, at which point the Hubble parameter is given by
$H\sim\frac{m_{\rm heavy}F^{2}}{f_{i}^{2}}\,,$ (12)
where we choose $i\,(=1,2)$ such that the strings associated with $\phi_{i}$ bound the DWs. Assuming $f_{1}\ll f_{2}$ for instance, $F\simeq f_{1}/n_{1}$ is much smaller than $f_{2}$. If $|n_{1}|=1$ and $|n_{2}|>1$, the strings associated with $\phi_{2}$ dominate the dynamics of the string-wall network, resulting in a delayed bundle formation by a factor of $(f_{2}/f_{1})^{2}$. Since the strings continue to dominate the string-wall network, the delayed formation does not lead to the DW problem. See also Ref. Eto:2023aqr for the case of hierarchical decay constants.
Figure 2: The snapshots of the energy density at the time $t=300/\sqrt{\lambda}\eta$ for $(n_{1},n_{2})=(1,2)$ (upper panel) and $(3,2)$ (lower panel) are shown. The energy density is normalized by $\lambda\eta^{4}$. In the upper panel, several string bundles, each composed of three strings, are formed. In contrast, a string-wall network is formed in the lower panel. In both panels, one side of the box is about the size of two Hubble horizons.
@@ III.2.2 Post-post scenario
On the other hand, if $D=0$, the $U(1)_{L}$ symmetry remains unbroken, and there is no DW attached to the string bundle. Thus, the string bundles without DWs simply follow the scaling solution. Let us mention here a potential difference from the simple single-axion case. Since the string bundles have an internal structure, it is nontrivial whether they evolve exactly in the same way as ordinary axion strings. For instance, we can consider a twisted loop formed by a string bundle. If the twist of the loop is preserved, the contraction of the loop will stop at a certain radius. Then, such twisted bundle loops are stable and may contribute to dark matter. However, we find that a string of one axion can easily pass through strings of the other axion in a lattice simulation. As a result, the twisted loop is untwisted and decays similarly to a loop of a single string. One way to prevent such untwisting is to introduce a repulsive interaction between the strings. Further exploration of such twisted loops will be presented in a forthcoming paper. Lee:preparation
@@ III.4 Summary of classification
$\geq 1$ $1$ 00 $\theta_{L}$ string bundles $\theta_{L}$ string bundles Figs. 2 and 8 -
@@ IV Cosmological implications
Alternatively, the string-wall network can collapse due to a transient bias that arises before the axions settle to the global minimum of the potential. This transient bias may either directly collapse the network or indirectly destabilize the network by introducing a population bias. As a result, the timescale of the string-wall network collapse can vary over a wide range. Here, we briefly discuss the condition for the transient bias to be effective. For simplicity, we assume that all the DW numbers are of $\mathcal{O}(1)$ and neglect them in the following. We can estimate how the potential bias affects the DW evolution by comparing the energy of the DWs and potential bias in each Hubble volume. The former is $\sim\sigma H^{-2}$ while the latter is $\sim\Lambda^{\prime 4}H^{-3}$. Here, the tension of the DW of $V_{1}$, $\sigma$, is given by $\sim\Lambda^{2}F$. Thus, the transient bias is effective if it is present when $H\lesssim\Lambda^{\prime 4}/(\Lambda^{2}F)$. Since $\phi_{\mathrm{light}}$ starts to oscillate when $H\simeq m_{\mathrm{light}}\sim\Lambda^{\prime 2}/\sqrt{f_{1}^{2}+f_{2}^{2}}$, the transient bias is effective if $\Lambda^{\prime 2}/\sqrt{f_{1}^{2}+f_{2}^{2}}\lesssim\Lambda^{\prime 4}/(\Lambda^{2}F)$, i.e., $f_{1}f_{2}/(f_{1}^{2}+f_{2}^{2})\lesssim\Lambda^{\prime 2}/\Lambda^{2}$. To realize such a situation, we need a hierarchy between $f_{1}$ and $f_{2}$ comparable to or more significant than that between $\Lambda^{2}$ and $\Lambda^{\prime 2}$.
@@
Thus, the key features of the QCD axion model with multiple axions are the general existence of high-tension DW networks and the generation of gravitational waves accompanying their collapse. This is precisely what was pointed out in the context of the clockwork QCD axion in Ref. Higaki:2016jjh .555 Ref. Higaki:2016jjh considered post-inflationary conditions and stated that if $N_{\rm DW}=1$, DWs would generally collapse if there were $N_{\rm axion}$ potential terms for $N_{\rm axion}$ axions. However, we have pointed out that this part is not correct, as the existence of induced walls was overlooked. Here, we have seen that this holds true even in the case of multiple axions, particularly for two axions.
@@ V Summary
We introduced two hierarchical axion potential terms and examined the evolution of topological defects based on their DW numbers. In the post-post scenario, configurations of string bundles with heavy axions confined within them exist but form effectively only under specific initial conditions and DW numbers. More generally, we observe stable networks with high-tension DWs caused by heavy axions. This challenges the common assumption that heavy axions can be integrated out to simplify the system to a single-axion model, especially in the presence of cosmic strings. In the pre-post scenario, the formation of such string bundles is impossible due to one set of strings being diluted away by inflation, inevitably leading to DW networks. The ubiquity of these DW networks in multi-PQ scalar scenarios is a key argument of this paper, generalizing the observation made by Higaki:2016jjh in the context of clockwork axions.
@@ Acknowledgments
Appendix A String bundle formation
Figure 8: The snapshots of strings and DWs at the time when $H^{-1}=483.8/\sqrt{\lambda}\eta$ for $(n_{1},n_{2})=(1,2)$ are shown. The blue (red) lines correspond to 1-strings (2-strings) and the semitransparent white surfaces correspond to heavy axion DWs. One side of the box is the size of two Hubble horizons.
We have discussed the heavy axion DW formation and the subsequent string bundle formation when $|n_{1}|=1$ and $|n_{2}|={\cal O}(1)$ (or vice versa) in the post-post scenario in Sec. III.1.1. Once the DWs are formed, they start to dominate the dynamics of strings in their scaling regime, and they will contract within ${\cal O}(1)$ Hubble time producing string bundles. This has been confirmed with 2D simulations (see Fig. 2).
Here, we show the numerical results from 3D lattice simulations to confirm that our expectations for the string bundle formation hold when three-dimensional structures are included, and thus the late-time dynamics of defects converges to that of a single-axion scenario.
We set $n_{1}=1$, $n_{2}=2$, $f_{1}=f_{2}\equiv\sqrt{2}\eta$, and $\lambda_{1}=\lambda_{2}\equiv\lambda$. The mass of $\phi_{\rm heavy}$ is set to be $m_{\rm heavy}^{2}=0.2\lambda\eta^{2}$. We start the simulation from the time when $H^{-1}=1/\sqrt{\lambda}\eta$ with random field values for each site. Fig. 8 shows a snapshot of the spatial distribution of strings and DWs in the 3D lattice. Note that the snapshot contains $2^{3}$ Hubble horizons. Most neighboring strings are aligned, and all the visibly remaining DWs have sizes smaller than the Hubble horizon, so we expect the string bundles to form soon. The shrinking of the DWs can be seen in Fig. 9, which plots the time evolution of the area parameter of the DWs, ${\cal A}_{3}$, normalized by $n_{2}$. The area parameter indicates the number of DWs per one Hubble horizon patch, and the detail of the definition is the same as in Ref. Kitajima:2023kzu ; we count the lattice sites where the field value $\phi_{\rm heavy}$ crosses the potential hilltop.
Here, we comment on the junction structures of $\phi_{\rm light}$ string networks, which can be seen in the snapshot. The junctions would be a possible deviation from the dynamics of the single-axion string, if they continued to exist. However, these junctions are considered to be unstable, because they disappear once pair annihilations or interconnections of strings occur. Even in the presence of interactions between the strings, they easily pass through each other due to the motion induced by the string tension. Thus, the junctions are only transient objects, and do not affect the overall time evolution of the string network.
Figure 9: The time evolution of DW area parameter ${\cal A}_{3}$ normalized by $n_{2}$. Note that the strings are formed and start to follow their scaling solutions at $2t=H^{-1}\simeq 60/\sqrt{\lambda}\eta$, and two scaling solutions of the strings become different each other due to the confinement of the DWs after $H^{-1}\simeq 90/\sqrt{\lambda}\eta$. The area parameter decreases and asymptotes to zero as the string bundles are being formed.